# GPT-4V with Emotion: A Zero-shot Benchmark for Generalized Emotion Recognition

### 03 Feb 2024 v2

@@ 1 Def Main($\mathcal{D}$):
3             Func($\mathcal{B}_{i}$);
@@ 8 Def Func($\mathcal{B}_{i}$):
9       for $t=1,\cdots,T_{\max}$ do
10             Call GPT-4V for $\mathcal{B}_{i}$ and get outputs $\mathcal{O}_{i}$; if passed checks then
11                   Store the outputs $\mathcal{O}_{i}$; return
@@
15      if $|\mathcal{B}_{i}|\geq 2$ then
@@
4.1 Batch-wise Calling Strategy
The current GPT-4V API has three limitations on requests: tokens per minute (TPM), requests per minute (RPM), and requests per day (RPD), which introduces additional requirements to our prompt design. To meet the constraints of RPM and RPD, we follow previous work wu2023gpt4vis and adopt batch-wise inputs. That is, we feed multiple samples into GPT-4V and request it to generate the response for each sample. However, a large batch size may result in the total number of tokens exceeding the TPM limit. Additionally, it increases task difficulty and may lead to incorrect outputs. For example, if we input 30 samples in a batch, we might only receive 28 predictions. Therefore, we set the batch size to 20 for image-level inputs and 6 for video-level inputs, aiming to satisfy three API limitations (i.e., TPM, RPM, and RPD) simultaneously. Prompts for different tasks can be found in Table 2.
4.2 Recursive Calling Strategy
During the evaluation, GER tasks tend to trigger security checks in GPT-4V. This is attributed to the inclusion of visual sentiment analysis and human emotion recognition. The former task contains violent and bloody images. In the latter task, human identities are also considered as sensitive information. To reduce these errors, we require GPT-4V to ignore speaker identity in the prompt (see Table 2), but it still triggers security errors. Interestingly, these errors sometimes occur randomly. For example, although all images are human-centric, some pass security checks while others fail. Alternatively, a sample may initially fail the check but pass upon retry. Therefore, we make multiple calls to the same batch to reduce rejected cases.
Meanwhile, we notice that a batch input may fail security checks, but splitting it into smaller parts can sometimes pass the checks. Therefore, for consistently failing batches, we split them into two smaller mini-batches and feed these mini-batches into GPT-4V. This operation is repeated until further splitting is no longer feasible. We call this strategy the “recursive calling strategy” and its pseudo-code is summarized in Algorithm 1.
@@ 5 Results and Discussion
In this section, we first show the performance of the heuristics baselines, supervised systems, and GPT-4V on GER tasks. Then, we reveal the multimodal fusion and temporal modeling capabilities of GPT-4V. Following that, we visualize confusion matrices and conduct an in-depth class-wise performance analysis. Next, we analyze the system stability and explore the color space robustness. Finally, we visualize the samples that are rejected in the security check and erroneously predicted.
@@
Model CMU-MOSI CH-SIMS MER-MULTI
MFM tsai2018learning 78.34 87.14 70.48
MISA hazarika2020misa 79.08 89.82 82.42
MFN zadeh2018memory 77.78 87.00 77.40
MMIM han2021improving 79.38 88.68 81.01
TFN zadeh2017tensor 79.63 90.56 82.52
MulT tsai2019multimodal 81.41 91.07 82.94
Random 51.33 51.23 17.87
Majority 42.37 50.47 10.40
GPT-4V (3-frm) 80.43 81.24 65.39
@@ 5.1 Main Results
In this section, we report the performance of different methods on five GER tasks. In Tables 3$\sim$7, we include two additional heuristic baselines: random guessing and majority guessing. For the former, we randomly select a label from the candidate categories as the prediction. For the latter, we choose the majority label as the prediction. For both baselines, we run experiments 10 times and report the average results.
Table 3 shows the results of visual sentiment analysis. From this table, we observe that GPT-4V outperforms supervised systems on most datasets. This superior performance can be attributed to its strong understanding of image content, coupled with its reasoning capabilities, allowing GPT-4V to accurately infer the emotional states evoked by the images. But for micro-emotion recognition (see Table 4), GPT-4V exhibits poor performance, even worse than heuristic baselines. These results suggest that GPT-4V is designed for general-purpose domains (i.e., emotions that can be perceived by normal people). It is not well-suited for tasks that require professional knowledge.
Tables 5$\sim$7 show the performance of facial emotion recognition, dynamic facial emotion recognition, and multimodal emotion recognition, respectively. The latter two tasks identify emotions in videos. Since GPT-4V does not support video input, we uniformly sample frames from the video and then feed these frames into GPT-4V sequentially. To reduce GPT-4V call costs, we sample up to three frames per video. Despite a performance gap still existing between GPT-4V and supervised systems, it is worth noting that GPT-4V significantly outperforms the heuristic baselines. These results demonstrate its potential in emotion recognition. In Table 7, GPT-4V performs well on CMU-MOSI but relatively poorly on MER-MULTI. This discrepancy can be attributed to the fact that acoustic information is more important in MER-MULTI compared to CMU-MOSI lian2024merbench . Since GPT-4V does not support audio input, MER-MULTI loses more information, leading to limited performance.
5.2 Multimodal Understanding Ability
In this section, we evaluate the multimodal understanding capabilities of GPT-4V. Among all tasks, only multimodal emotion recognition provides multimodal information. Therefore, we conduct experiments on this task. Table 8 reports both unimodal and multimodal results. For CH-SIMS and MER-MULTI, we observe that multimodal results outperform unimodal results, showcasing GPT-4V’s ability to integrate and leverage multi-modalities. But for CMU-MOSI, we note a slight decrease in multimodal results compared to unimodal results. This can be attributed to CMU-MOSI primarily relying on lexical information to convey emotions, and the incorporation of visual information may introduce interference lian2024merbench .
5.3 Temporal Understanding Ability
@@
Table 8: Analysis of multimodal fusion ability. We provide both unimodal and multimodal results for the multimodal emotion recognition task.
Modality CMU-MOSI CH-SIMS MER-MULTI
Text 82.32 70.07 34.57
Video 51.17 76.13 46.23
Multi 80.43 81.24 65.39
@@
(b) Twitter I(4)
(c) Twitter I(3)
(d) Twitter II
(e) Abstract
(f) ArtPhoto
@@
5.4 Confusion Matrix
In Figure 3, we visualize confusion matrices and analyze the performance of GPT-4V across various emotion categories.
For visual sentiment analysis, GPT-4V generally performs well in both positive and negative classes, showcasing its robust visual understanding capabilities. However, it exhibits lower performance on the Abstract dataset compared to other datasets. This may be attributed to the fact that GPT-4V is mainly trained on natural images. The domain gap between abstract and natural images leads to increased difficulty in emotion recognition.
For micro-expression recognition, GPT-4V prefers tense and repression. Compared to other labels (e.g., surprise and disgust), tense and repression are less noticeable and closer to neutral. These results indicate that GPT-4V tends to view micro-expressions as neutral, causing it to perform poorly on this task.
For facial expression recognition, GPT-4V tends to exhibit lower accuracy on contempt, disgust, and fear, while achieving higher accuracy on happiness. This phenomenon is related to the clarity of emotion definition. Since contempt, disgust, and fear have less consensus between annotations, they are inherently more difficult to recognize mollahosseini2017affectnet ; picard2001toward .
For dynamic facial emotion recognition, most samples tend to be predicted as neutral, which is more obvious in the FERV39 dataset. We attribute this phenomenon to our sampling operation. For each video, we sample up to three frames. If there is no emotion-related information within these frames, GPT-4V will predict the emotion as neutral.
For multimodal emotion recognition, GPT-4V performs well on worry and happiness. For worry, it has a strong reliance on lexical information. Differently, happiness is mainly expressed through facial expressions lian2023mer . To achieve good performance, GPT-4V should use different predominant modalities for distinct emotions, showcasing its multimodal fusion capabilities.
(a) identical predictions
(b) test accuracy
5.5 System Stability
In this section, we evaluate the system stability of GPT-4V. Experiments are conducted on SFEW 2.0, using GPT-4V to predict each sample 10 times. Figure 4 calculates the frequency with identical predictions and reports the test accuracy for different runs. As shown in Figure 4a, although most samples exhibit identical predictions across 10 runs, there are some samples with different predictions, leading to fluctuations in the test accuracy. In Figure 4b, we observe that the maximum performance gap between the best and worst runs is 4.60%. Hence, GPT-4V exhibits some instability. We recommend that subsequent researchers evaluate GPT-4V in multiple runs and consider the majority predictions as the final results.
@@ 5.6 Robustness to Color Space
As illustrated in Figure 2, there are differences in the color spaces across distinct datasets. In this section, we evaluate the robustness of GPT-4V to color space changes. We conduct experiments on RAF-DB and convert all RGB images to grayscale. We observe that GPT-4V achieves nearly identical performance on grayscale images (74.28%) and RGB images (75.81%). These results demonstrate its color space robustness.
Figure 5 further visualizes the prediction consistency on grayscale and RGB images. We observe that the consistency rate is around 80% for most emotion categories, except for fear and disgust. We further compare Figure 5 with the confusion matrix of RAF-DB (see Figure 3m). We observe significant similarities between these images, where fear and disgust also perform poorly in emotion recognition. Therefore, a reasonable conjecture emerges: poorly-performing emotion categories in GPT-4V may exhibit greater randomness across different runs.
@@ 5.7 Security Check
GER includes visual sentiment analysis (containing images with violent and bloody content) and human emotion recognition (where human identity is considered sensitive information in GPT-4V). Therefore, it often triggers the security check. Figure 6 shows passed and rejected examples. We observe that all rejected examples contain humans or violent scenes. But there are some cases where examples containing such information can still pass checks. This observation suggests a certain degree of instability in the current security check used by GPT-4V.
@@ 5.8 Case Study
In Figure 7, we visualize examples where GPT-4V makes incorrect predictions. More examples can be found in Figures 8$\sim$12 in the Appendix. For visual sentiment analysis (see Twitter I), we notice that GPT-4V can generate reasonable predictions. The main reason for the errors is the inherent multiple meanings of images. Take the image on the left as an example. One interpretation could emphasize that the sun rises, breaking through dark clouds and bathing the trees in sunlight, evoking positive emotions. Conversely, another explanation could emphasize that the sun is obstructed by dark clouds, triggering negative emotions. The same phenomenon can also be found in the picture on the right. On the one hand, there is a cute cat, but on the other hand, the cat is locked in a cage.
Figure 6: Passed and rejected examples by GPT-4V.
@@
For micro-expression recognition, we observe that her emotion is not obvious (see CASME in Figure 7). Without specialized knowledge, it is difficult to identify her emotion as disgust and more likely to be predicted as neutral. For CASME, there are four candidate labels: tense, disgust, repression, and surprise. Among them, tense is closer to neutral. Consequently, GPT-4V incorrectly predicts her emotion as tense. In the future, we recommend that researchers provide clearer definitions of micro-expressions in the prompt design. For example, offer typical clues to distinguish these emotions. We believe this strategy can enhance the performance of GPT-4V on this task.
@@
In summary, GPT-4V provides relatively reasonable predictions. These errors are primarily caused by the inherent multiple meanings of images or the limited capability to identify small objects. Additionally, micro-expressions are different from regular expressions, and GPT-4V encounters challenges in recognizing them without specialized knowledge. Furthermore, its multimodal fusion ability may face challenges when different modalities express distinct emotions.

### 18 Mar 2024 v3

@@ 1 Def Main($\mathcal{D}$):
3             # batch-wise calling Func($\mathcal{B}_{i}$);
@@ 8 Def Func($\mathcal{B}_{i}$):
9       # repeated calling for $t=1,\cdots,T_{\max}$ do
10             Call GPT-4V for $\mathcal{B}_{i}$ and get response $\mathcal{O}_{i}$; if meet requirements then
11                   Store the response $\mathcal{O}_{i}$; return
@@
15      # recursive calling if $|\mathcal{B}_{i}|\geq 2$ then
@@
4 GPT-4V Calling Strategy
This paper evaluates the performance of the GPT-4V API, “gpt-4-vision-preview”. GER tasks involve diverse modalities, such as images, text, video, and audio. However, GPT-4V only supports image and text inputs. To process video data, we sample the video and convert it into multiple images. To process audio data, we try converting the audio to a mel-spectrogram, which is a typical visual representation of the frequency content of an audio signal over time, with the frequency axis scaled according to the mel scale hermansky1990perceptual . However, GPT-4V fails to generate proper responses for the mel-spectrogram. Therefore, the primary focus of this paper revolves around images, text, and video. In this section, we design a calling strategy specifically for GER tasks, containing batch-wise, repeated, and recursive calling. Pseudo-code is summarized in Algorithm 1.
Batch-wise Calling. GPT-4V API has three limitations on requests: tokens per minute (TPM), requests per minute (RPM), and requests per day (RPD), which introduces additional requirements to our prompt design. To meet the constraints of RPM and RPD, we follow previous work wu2023gpt4vis and adopt batch-wise inputs. That is, we feed multiple samples into GPT-4V and generate all results in one request. However, a large batch size may result in the total number of tokens exceeding the TPM limitation. Additionally, it increases task difficulty and may lead to incorrect responses. For example, if we input 100 samples in a request, GPT-4V might only generate predictions for 96 samples. Therefore, we set the batch size to 20 for image-level inputs and 6 for video-level inputs, aiming to satisfy TPM, RPM, and RPD limitations simultaneously. Prompts for different tasks can be found in Table 2.
Repeated Calling. GER tasks often trigger security checks, causing GPT-4V to refuse to provide a response. This is attributed to the inclusion of visual sentiment analysis and human emotion recognition. The former task contains violent and bloody images. In the latter task, human identities are also considered as sensitive information. To reduce rejected cases, we require GPT-4V to ignore speaker identity (see Table 2), but it still triggers security errors. Interestingly, these errors sometimes occur randomly. For example, although all images are human-centric, some pass security checks while others fail. Alternatively, a sample may initially fail the check but pass upon retry. Therefore, we make multiple calls to the rejected batches.
Recursive Calling. During the evaluation, we notice that a batch-wise input may fail security checks, but splitting it into smaller parts can sometimes pass checks. Therefore, for the consistently failing batch, we split it into two smaller mini-batches and feed them into GPT-4V separately. This operation is repeated until further splitting is no longer feasible.
Combination. Our strategy combines batch-wise, repeated, and recursive calling. More details can be found in Algorithm 1. A qualified response $\mathcal{O}$ should satisfy two conditions. First, it should not trigger the security check. Second, it should contain the correct number of prediction results. Figures 8$\sim$13 provide examples of qualified responses for different GER tasks.
@@ 5 Results and Discussion
This section presents and discusses results from three levels: dataset-wise, class-wise, and sample-wise. Specifically, we first present the results of heuristics baselines, supervised systems, and GPT-4V. Then, we reveal the temporal modeling and multimodal fusion capabilities of GPT-4V. Following that, we analyze the system’s stability and conduct an in-depth class-wise performance analysis. Next, we assess the system’s robustness against changes in prompt templates and color spaces. Finally, we visualize the rejected and erroneously predicted samples.
@@
Table 5: Performance comparison on tweet sentiment analysis.
Model MVSA-Single MVSA-Multiple
CNN-Multi cai2015convolutional 61.20 66.39
DNN-LR yu2016visual 61.42 67.86
MultiSentiNet xu2017multisentinet 69.84 68.86
CoMN xu2018co 70.51 69.92
MVAN yang2020image 72.98 72.36
ITIN zhu2022multimodal 75.19 73.52
Random 33.96 33.46
Majority 55.68 66.65
GPT-4V 61.25 66.82
Model CMU-MOSI CH-SIMS MER-MULTI
MFM tsai2018learning 78.34 87.14 70.48
MISA hazarika2020misa 79.08 89.82 82.42
MFN zadeh2018memory 77.78 87.00 77.40
MMIM han2021improving 79.38 88.68 81.01
TFN zadeh2017tensor 79.63 90.56 82.52
MulT tsai2019multimodal 81.41 91.07 82.94
Random 51.33 51.23 17.87
Majority 42.37 50.47 10.40
GPT-4V 80.43 81.24 65.39
@@ 5.1 Main Results
Besides GPT-4V, we introduce two heuristic baselines: random guessing and majority guessing. For the former, we randomly select a label from candidate labels. For the latter, we choose the majority label as the prediction. For both baselines, we run experiments 10 times and report average results.
Table 3 shows the results of visual sentiment analysis. We observe that GPT-4V outperforms supervised systems on most datasets. This superior performance can be attributed to its strong visual understanding ability, coupled with its reasoning capability, allowing GPT-4V to accurately infer the emotional states evoked by images. But for micro-expression recognition (see Table 4), GPT-4V exhibits poor performance, even worse than heuristic baselines. These results suggest that GPT-4V is designed for general-purpose domains (i.e., emotions that can be perceived by normal people). It is not well-suited for tasks that require professional knowledge like micro-expressions.
Tables 5$\sim$8 present the results of tweet sentiment analysis, multimodal emotion recognition, facial emotion recognition, and dynamic facial emotion recognition, respectively. To process video data, we uniformly sample frames and feed these frames into GPT-4V sequentially. To reduce call costs, we sample up to three frames per video. Despite a performance gap still existing between GPT-4V and supervised systems, it is worth noting that GPT-4V can significantly outperform heuristic baselines, demonstrating its potential in emotion recognition.
In Table 6, we observe a larger performance gap between supervised systems and GPT-4V on MER-MULTI compared to CMU-MOSI and CH-SIMS. This can be attributed to the fact that supervised systems incorporate acoustic information, while GPT-4V does not support audio input. Given that audio is more crucial in MER-MULTI lian2024merbench , GPT-4V loses more information in this dataset, leading to limited performance.
5.2 Temporal Modeling Ability
@@
5.3 Multimodal Fusion Ability
This section evaluates GPT-4V’s multimodal fusion ability. Among all tasks, tweet emotion recognition and multimodal emotion recognition provide multimodal information. Hence, we conduct experiments on these tasks. Table 9 shows unimodal and multimodal results. In general, multimodal results outperform unimodal results, showcasing GPT-4V’s ability to integrate and exploit multi-modalities. But for CMU-MOSI, we notice a slight decrease in multimodal results compared to unimodal results. This can be attributed to CMU-MOSI mainly relying on lexical content to convey emotions lian2024merbench , and the incorporation of visual clues may introduce interference.
Table 9: Multimodal fusion ability. We report unimodal and multimodal results for tweet sentiment analysis (TSA) and multimodal emotion recognition (MER). Here, TSA takes text-image inputs and MER takes text-video inputs.
Task Dataset Text Image/Video Fusion
TSA MVSA-Single 57.65 58.68 61.25
MVSA-Multiple 62.71 63.35 66.82
MER CH-SIMS 70.07 76.13 81.24
CMU-MOSI 82.32 51.17 80.43
MER-MULTI 34.57 46.23 65.39
5.4 System Stability
This section evaluates GPT-4V’s prediction stability. During the experiments, we run GPT-4V 10 times for each sample in SFEW 2.0. Figure 3a shows the frequency of identical predictions. Specifically, we assume that GPT-4V predicts negative 8 times and negative 2 times. Therefore, it predicts the same label at most $c=8$ times. Then, we compute $c$ for all samples and calculate its frequency. In Figure 3b, we report the test accuracy for each run. We observe that although more than 50% of the samples exhibit the same result across 10 runs, there are also some samples with different prediction results across distinct runs, leading to fluctuations in test accuracy. Meanwhile, we observe a 4.60% performance gap between the best and worst runs. Therefore, GPT-4V exhibits certain instability. We recommend that subsequent researchers evaluate GPT-4V multiple times and use majority voting to get final results.
5.5 Class-wise Performance Analysis
In Figure 4, we visualize confusion matrices and perform a class-wise performance analysis. For visual sentiment analysis, GPT-4V generally performs well in positive and negative classes, showcasing its visual understanding ability. However, it exhibits lower results on Abstract compared to other datasets. This may be attributed to the fact that GPT-4V is mainly trained on natural images. The domain gap between abstract and natural images leads to increased difficulty and limited performance. For tweet sentiment analysis, GPT-4V performs relatively poorly in identifying neutral sentiment and often misclassifies it as positive or negative. This may be due to GPT-4V’s ability to detect subtle emotional cues in image-text pairs.
(a) identical predictions
(b) test accuracy
@@ 5.5 Class-wise Performance Analysis
(b) Twitter II
(c) Abstract
(d) ArtPhoto
(e) MVSA-Single
(f) MVSA-Multiple
@@
Table 10: Robustness to prompt template changes. We conduct experiments on the facial emotion recognition task.
ID
Prompt Template
#1
#2
We provide 20 images. For each image, please sort the provided categories from high to low according to the top 5 similarity with the input image. Here are the optional categories: {Candidate labels}. Please ignore the speaker’s identity and focus on the facial expression. The output format should be {’name’:, ’result’:} for each image.
#3
Please play the role of a facial expression classification expert. We provide 20 images. For each image, please select the most likely category according to the correlation with the input image. Here are the optional categories: {Candidate labels}. Please ignore the speaker’s identity and focus on the facial expression. The output format should be {’name’:, ’result’:} for each image.
For micro-expression recognition, GPT-4V prefers tense and repression. Compared to other labels (e.g., surprise and disgust), tense and repression are less noticeable and closer to neutral. These results indicate that GPT-4V tends to view micro-expressions as neutral, causing it to perform poorly on this task. For facial expression recognition, GPT-4V tends to exhibit lower accuracy on contempt, disgust, and fear, while achieving higher accuracy on happiness. This phenomenon is related to the clarity of emotion definition. Since contempt, disgust, and fear have less consensus between different annotations, they are inherently more difficult to recognize mollahosseini2017affectnet ; picard2001toward .
For dynamic facial emotion recognition, most samples tend to be predicted as neutral, which is more obvious in FERV39. We attribute this phenomenon to our sampling operation. For each video, we sample up to three frames. GPT-4V tends to predict the emotion as neutral if there is no emotion-related information within these frames. For multimodal emotion recognition, GPT-4V performs well on worry and happiness. For worry, it has a strong reliance on lexical information. Differently, happiness is mainly expressed through facial expressions lian2023mer . To achieve good performance, GPT-4V should use different predominant modalities for distinct emotions, showcasing its multimodal fusion capabilities.
5.6 Robustness to Template Change
Previous works have pointed out that pre-trained language models (such as BERT and RoBERTa) are sensitive to prompt templates, with different templates yield varying results mao2022biases . This raises the question: is GPT-4V also sensitive to template changes? In this section, we conduct experiments on SFEW 2.0 and compare the performance of different templates. In Table 10, Template#1 is our default template. Compared with Template#1, Template#2 deletes “Please play the role of a facial expression classification expert” and Template#3 changes the selection of the top 5 classes into the selection of the most likely class. Experimental results demonstrate that Template#2 (57.93%) and Template#3 (54.25%) can achieve similar performance compared to Template#1 (57.24%). Therefore, we can conclude that GPT-4V is robust to template changes.
@@ 5.7 Robustness to Color Space
This section evaluates the robustness of GPT-4V to color space changes. We conduct experiments on RAF-DB and convert RGB images into grayscale images. We observe that GPT-4V achieves nearly identical results on grayscale (74.28%) and RGB images (75.81%). These results demonstrate its robustness to color space changes. In Figure 5, we further analyze the class-wise prediction consistency. We observe that the consistency rate is around 80% for most classes except for fear and disgust. We further compare Figure 5 with the confusion matrix in Figure 4m. We observe similarities between these figures, where fear and disgust also perform poorly in emotion recognition. Therefore, we can conclude that poorly-performing categories in GPT-4V exhibit greater inconsistency in different color spaces. This may be attributed to poorly-performing categories with greater randomness across different runs.
@@ 5.8 Security Check
GER includes visual sentiment analysis (containing images with violent and bloody content) and human emotion recognition (where human identity is considered sensitive information in GPT-4V). Therefore, it often triggers the security check. Figure 6 visualizes passed and rejected examples. In general, rejected examples contain humans or violent scenes. But there are some cases where examples containing such information can still pass checks. This observation suggests a certain degree of instability in the current security check used by GPT-4V.
@@ 5.9 Case Study
In Figure 7, we visualize examples where GPT-4V makes incorrect predictions. More examples can be found in Figures 8$\sim$13 in the Appendix. For visual sentiment analysis (see Twitter I), we notice that GPT-4V can generate reasonable predictions. The main reason for these errors is the inherent multiple meanings of images. Take the image on the left as an example. One interpretation could emphasize that the sun rises, breaking through dark clouds and bathing the trees in sunlight, evoking positive emotions. Conversely, another explanation could emphasize that the sun is obstructed by dark clouds, triggering negative emotions. The same phenomenon can also be found in the picture on the right. On the one hand, there is a cute cat, but on the other hand, the cat is locked in a cage.
@@
For tweet sentiment analysis (see MVSA-Multiple), the image shows a hat on wood, accompanied by descriptive text about the hat. Although the label is neutral, GPT-4V predicts a positive sentiment. We speculate that GPT-4V may interpret this image as depicting a comfortable environment and a nice hat, thereby eliciting a positive response.
For micro-expression recognition (see CASME), we observe that her emotion is not obvious. Without specialized knowledge, it is difficult to identify her emotion as disgust and more likely to be predicted as neutral. For CASME, there are four candidate labels: tense, disgust, repression, and surprise. Among them, tense is closer to neutral. Consequently, GPT-4V incorrectly predicts her emotion as tense. In the future, we recommend that researchers provide clearer definitions of micro-expressions in the prompt. For example, offer typical clues to distinguish these micro-expressions. We believe this strategy can enhance the performance of GPT-4V on this task.
@@
In summary, GPT-4V can provide relatively reasonable responses. These errors are mainly caused by the inherent multiple meanings of images or the limited capability to understand the details. Meanwhile, micro-expressions are different from regular emotions and GPT-4V faces challenges in recognizing them. Furthermore, its multimodal fusion ability may face challenges when different modalities express distinct emotions.
6 Limitations and Future Directions
In this section, we summarize the main challenges and limitations of GPT-4V as revealed by the evaluation results. Additionally, we discuss potential directions for future research.
Performance. In GER tasks, GPT-4V outperforms heuristic baselines but still lags behind supervised systems. This may be attributed to the vague definition of emotion lian2023explainable . In the future, we plan to use few-shot prompts to enhance task clarity and GPT-4V’s emotion understanding ability.
Support Modality. GPT-4V does not support audio. However, audio can convey emotions through prosodic information such as pitch, duration, and intensity. To better recognize emotions, GPT-4V should support audio in the future. Alternatively, we can use outside models to convert audio into lexical descriptions and use them as additional input for GPT-4V.
Domain Gap. GPT-4V is designed for general-purpose domains and performs poorly on micro-expressions that require specialized knowledge. For future work, we recommend that subsequent researchers can attempt to provide clearer definitions of micro-expressions in the prompt design.
System Stability. GPT-4V exhibits certain instability in prediction results, i.e., it may predict different labels for a sample. Therefore, GPT-4V should improve its system stability in the future. Alternatively, subsequent works should evaluate GPT-4V multiple times and use majority voting to get final results.
Security Check Stability. GPT-4V demonstrates certain instability in security checks. Specifically, a sample may initially fail the check but pass after a retry. To solve this problem, we design a more practical calling strategy in Section 4. In the future, GPT-4V should enhance its security check stability.
Despite the above limitations, GPT-4V has demonstrated promising results in emotion recognition and holds application potential in practical scenarios such as social media analysis, online education, and customer service. Different from previous solutions that required task-specific models tailored to each scenario, GPT-4V can serve as a universal model capable of minimizing application costs.